An object is kept at x = 0, a force on the object is F (constant) from t =0. Friction coefficient vary as μ=cx. Then the distance travelled by the block is

Grade:11

1 Answers

Vikas TU
7 years ago
Distance travelled by the block =>
Net Force on block = > 
F – uMg = Ma
a =  acceleration of block.
F = M[cgx + vdv/dx]
vdv = Fdx/M  –  cgxdx
integerating both sides,
v^2/2 = Fx/M -cgx^2/2
v = root(2xF/M – cgx^2)
 
v would be zero at last.
hence, 2F/M – cgx = 0
or x = 2F/cgM is the distance covered by the block.